Consider the following general demand and supply functions: 𝑄𝐷 = 1800 βˆ’ 8𝑃 + 2𝑀 + 16𝑃𝑅 𝑄𝑆 = 50 + 12𝑃 βˆ’ 20𝑃𝐡 + 19𝐹 where QD and QS are quantities demanded and supplied, respectively, P is price of the good, M is household income, PR is price of a demand-related good, PB is price of a supply-related good, and F is number of firms in the industry


1
Expert's answer
2022-02-13T11:50:16-0500

In equilibrium Qd = Qs, so:

1800 βˆ’ 8𝑃 + 2𝑀 + 16𝑃𝑅 = 50 + 12𝑃 βˆ’ 20𝑃𝐡 + 19𝐹,

20P = 1750 + 2M + 16PR + 20PB - 19F,

P = 87.5 + 0.1M + 0.8PR + PB - 0.95F.